# Exploring Model Kinship for Merging Large Language Models

### v2

@@ 4.4 Discussion
Considering all results that we observed, this analysis provides two main inspirations for the application of model kinship:
  • High kinship merges may lead to performance stagnation, similar to biological inbreeding.
  • Low kinship merges involve risk, but can lead to greater gains, potentially enabling escape from local optima caused by greedy strategy.
Greedy Strategy + Model Kinship
Model Avg. Gain Kinship Model Avg. Gain Kinship
MetaMath 63.72 / / MetaMath 63.72 / /
Instruct 61.82 / / Instruct 61.82 / /
Open-chat 66.28 / / Open-chat 66.28 / /
model-1-1 62.17 -0.6 0.01 model-1-1 62.17 -0.6 0.01
model-1-2 64.02 -0.03 -0.02 model-1-2 64.02 -0.03 -0.02
model-1-3 66.84 +1.84 0.05 model-1-3 66.84 +1.84 0.05
model-2-1 68.72 +2.16 0.93 model-2-1 68.72 +2.16 0.93
model-2-2 61.47 -3.96 0.57 model-2-2 61.47 -3.96 0.57
model-2-3 61.32 -3.83 0.58 model-2-3 61.32 -3.83 0.58
model-3-1 68.59 +1.09 0.95 model-3-2 67.74 +1.09 0.93
model-3-2 67.74 -0.04 0.93 model-3-3 69.06 +0.74 0.24
- - - model-3-4 68.61 +1.13 0.32
model-4-1 68.51 -0.14 0.98 model-4-4 68.75 -0.14 0.54
model-4-2 68.04 -0.19 0.98 model-4-5 68.39 -0.27 0.66
model-4-3 68.53 +0.37 0.94 model-4-6 69.03 +0.15 0.52
- - - model-5-1 69.13 +0.04 0.65
- - - model-5-2 68.98 +0.07 0.65
- - - model-5-3 68.63 -0.37 0.98
@@ 5 Using Model Kinship to Improve Model Merging
Building on the insights from section 4, we explore how model kinship can be leveraged to improve the model merging process. Our main experiment centers on the Mistral-7B model, with detailed results presented in the main text. To further evaluate the generalization of our approach, we also conduct two supplementary experiments: one based on Llama-2 (see Appendix B) and another on a distinct task set to test robustness across different evaluation settings. Our results indicate that while greedy strategy focuses on short-term gains, it can lead to suboptimal outcomes. By integrating model kinship, we can help the strategy avoid local optima and gain further improvements.
@@ 5.1 Main Experiment Setup
For the main experiments, we follow the same settings as described in section 3, including the use of the three fine-tuned Mistral-7B variants and the evaluation on Winogrande, GSM8k, and TruthfulQA. We adopt the Top-$k$ Greedy Merging strategy as the baseline continual merging strategy.
@@ 5.2 Results and Discussion
Figure 5 (a) illustrates the improvements in top average task performance across merging generations. Table 3 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method. While both strategies achieve the multitask objective, the vanilla greedy strategy ceases to improve after Generation 2, plateauing at an average task performance of 68.72. In contrast, strategy utilizing model kinship escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
@@
Figure 5 (b) shows the main branch of the model family tree. To explore how low-kinship merges help escape local optima during saturation, we analyze weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3), with $v_{pre}$ (from Model-1-3 to Model-2-1) as baseline. Figure 6 shows that merging with the exploration model ($v_{2}$) induces significant, novel weight changes, while $v_{1}$ shows minimal change due to high similarity between openchat-3.5 and Model-2-1.
@@
Iterative merging is resource-intensive. In our main experiments, a greedy strategy saturated after 2/4 merges with no further gains. Looking back at community experiments, 5/14 merges in path 1 averaged only 0.57 improvement, and 3/12 merges in path 2 averaged 0.36. A high kinship score (PCC > 0.9) among top models may indicate convergence. Stopping merges early at high kinship generation saves  30% time with negligible performance loss.
@@ 6 Related work
Weight averaging, a widely used technique in model merging, traces its origins to Utans (1996). Since the 2010s, weight averaging has been widely applied in deep neural networks, notably for combining checkpoints to improve training stability and performance. (Nagarajan and Kolter, 2019; Tarvainen and Valpola, 2017; Izmailov et al., 2018; Li et al., 2023a; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023b; Smith and Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022). Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022). Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024). Ilharco et al. (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base. Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024), which prevents interference by randomly dropping weights. The Model Breadcrumbs (Davari and Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging. Merging models with different initializations requires additional considerations. Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh and Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024). While weight averaging cannot be applied to models with different architectures, it can still be used to enhance feasible fusion methods. Recent work like FuseChat (Wan et al., 2024b) integrates it with Knowledge Fusion (Wan et al., 2024a) to enable novel fusion approaches.
Recently, there have been some works exploring “model evolution”. Tellamekala et al. (2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multitask models. Labonne (2024) develop a tool to automatically merge models on HuggingFace. Akiba et al. (2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations.
@@ 7 Conclusion
We propose continual model merging as a framework for evolving large language models through strategic, iterative merges that yield consistent gains in generalization and task performance. To support this framework, we introduce model kinship, a metric that guides merge candidate selection and explains both performance gains and stagnation during merging. Building on this, we propose Top-$k$ Greedy Merging with Model Kinship, a strategy that uses kinship to escape local optima and achieve further improvements. Kinship also serves as an early stopping signal by detecting convergence and reducing redundant computation.
@@
However, there are several limitations to consider: a) The experiments in this study are conducted on models with two architectures, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu and Dao, 2023). Furthermore, scaling of tasks and candidate models requires further experimentation to understand the computational cost across various scenarios. b) The analysis is based on open source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias. c) Possible correlation metrics for model kinship have not been fully explored. Other metrics may perform better than those discussed in this paper. d) The effectiveness of model kinship is demonstrated through empirical evidence. However, a theoretical framework (such as the assumptions in Appendix 15) is needed to explain model evolution and model kinship more rigorously. e) Model kinship currently guides merging and enhances performance, but does not support sustained evolution. Future progress may require environmental feedback, reward models (Silver et al., 2021), as well as new architectures.
@@ A.1 Selecting the Evolution Path
  • Performance and Kinship Evaluation: These branches were analyzed for changes in merging performance, particularly focusing on shifts in multitask capabilities and model kinship metrics.
@@ D.1 Analysis of Model Kinship Change across Merging Stages
To determine whether the discovery of increasing model kinship in model evolution paths can be generalized to the entire model evolution process, we perform an analysis of the merging stages. Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of models at each stage. For this analysis, we randomly select 5 models from each merging stage, as delineated by boundaries identified in prior analysis - Saturation Stage ($\geq 0.75$), Improving Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
@@ Appendix E Analysis between Task Relatedness and Model Kinship
In the formulation of model kinship, we use the placeholder (,$\sim$($\cdot$,$\cdot$)∼ ( ⋅ , ⋅ ) as a similarity metric function to explore options that can effectively capture task-related differences. One such metric is cosine similarity, derived from the analysis in the task vector, which has been validated as effective for representing differences in single-task models through the cosine similarity of delta parameters (task vectors). In addition to cosine similarity, we also investigate the Pearson correlation coefficient and Euclidean distance.
However, we have not thoroughly evaluated the applicability of these metrics in the context of model evolution, particularly for merged models with multitask capabilities. To address this, we examine the relationship between the similarity metrics and task information in subsequent sections.
@@ Appendix F Optimization Assumption of Model Evolution
Figure 9: An intuitive illustration of how model evolution can fall into local optima due to a performance-prior strategy. It shows that Merged Model 2 may be overlooked, despite its potential for better multitask performance.
@@
Figure 11: An intuitive comparison between selective breeding and continual model merging. The left process demonstrates breeding a tall and frutful plant by selecting parents with the desired traits in an biological scenario. The right process shows developing a model with capabilities of coding and math through model evolution.

### v3

@@ 4.4 Discussion
Considering all results that we observed, this analysis provides two main insights for the application of model kinship:
  • High kinship merges may lead to performance stagnation, similar to biological inbreeding, where excessive similarity limits the ability to adjust.
  • Low kinship merges involve risk, but can result in greater gains, potentially enabling escape from local optima caused by the greedy strategy.
@@ 5 Using Model Kinship to Improve Model Merging
Building on the insights from section˜4, we explore how model kinship can be leveraged to improve the model merging process. Our main experiment centers on the Mistral-7B model, with detailed results presented in the main text. To further evaluate the generalization of our approach, we also conduct two supplementary experiments: one based on Llama-2 (see Appendix B) and another on a distinct task set to test robustness across different evaluation settings. Our results indicate that while greedy strategy focuses on short-term gains, it can lead to suboptimal outcomes. By integrating model kinship, we can help the strategy avoid local optima and gain further improvements.
@@ 5.1 Main Experiment Setup
For the main experiments, we follow the same settings as described in section˜3, including the use of the three fine-tuned Mistral-7B variants and the evaluation on Winogrande, GSM8k, and TruthfulQA. We adopt the Top-$k$ Greedy Merging strategy as the baseline iterative merging strategy [^2] We provide a Google Colab Notebook..
@@ 5.2 Results and Discussion
Figure 5 (a) illustrates the improvements in top average task performance across merging generations. Table 3 provides model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method. While both strategies achieve the multi-task objective, the vanilla greedy strategy ceases to improve after Generation 2, plateauing at an average task performance of 68.72. In contrast, strategy utilizing model kinship escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
@@
Figure 5 (b) shows the main branch of the model family tree. To explore how low-kinship merges help escape local optima during saturation, we analyze weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3), with $v_{pre}$ (from Model-1-3 to Model-2-1) as baseline. Figure 6 shows that merging with the exploration model ($v_{2}$) produces significant, novel weight changes, while $v_{1}$ shows minimal change due to the high similarity between openchat-3.5 and Model-2-1.
@@
Iterative model merging can be resource intensive. In our main experiments, a greedy strategy saturated after 2/4 merges with no further gains. Looking back at community experiments, 5/14 merges in path 1 averaged only 0.57 improvement, and 3/12 merges in path 2 averaged 0.36. A high kinship score (PCC > 0.9) among top models may indicate convergence. Stopping merges early at high kinship generation saves  30% time with negligible performance loss.
@@ 6 Related work
Weight averaging, a widely used technique in model merging, traces its origins to Utans (1996). Since the 2010s, weight averaging has been widely applied in deep neural networks, notably for combining checkpoints to improve training stability and performance. (Nagarajan and Kolter, 2019; Tarvainen and Valpola, 2017; Izmailov et al., 2018; Li et al., 2023a; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023b; Ilharco et al., 2022; Izmailov et al., 2018; Smith and Gashler, 2017), and parallel training of large language models (LLMs) (Li et al., 2022). Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022). Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024). Ilharco et al. (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base. Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024b), which prevents interference by randomly dropping weights. The Model Breadcrumbs (Davari and Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging. Merging models with different initializations requires additional considerations. Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh and Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024). While weight averaging cannot be applied to models with different architectures, it can still be used to enhance feasible fusion methods. Recent work like FuseChat (Wan et al., 2024b) integrates it with Knowledge Fusion (Wan et al., 2024a) to enable novel fusion approaches. Li et al. (2025) further demonstrate that model merging can stabilize training and serve as a low-cost simulator for annealed performance during pretraining, enabling checkpoint reuse and faster validation.
Recently, there have been some works exploring “model evolution”. Tellamekala et al. (2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multi-task models. Labonne (2024) develop a tool to automatically merge models on HuggingFace. Akiba et al. (2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations. Tang et al. (2025) present a continual model merging method that enables training-free projection-based merging of models as they arrive sequentially, improving efficiency and task retention without retraining.
@@ 7 Conclusion
We propose iterative model merging as a framework for evolving large language models through strategic, iterative merges that yield consistent gains in generalization and task performance. To support this framework, we introduce model kinship, a metric that guides merge candidate selection and explains both performance gains and stagnation during merging. Building on this, we propose Top-$k$ Greedy Merging with Model Kinship, a strategy that uses kinship to escape local optima and achieve further improvements. Kinship also serves as an early stopping signal by detecting convergence and reducing redundant computation.
@@
This study has several limitations: a) Our experiments are limited to two model architectures. It remains unclear whether our metric and method generalize to other architectures, such as Mamba (Gu and Dao, 2023). In addition, the scalability of tasks and candidate models requires further evaluation to assess computational costs in diverse settings. b) The analysis relies on open-source data from the Open Leaderboard. As this dataset is community-generated, it may be affected by noise or user bias. c) We have not fully explored alternative correlation metrics for measuring model kinship. Other metrics may yield stronger or more consistent performance. d) Our demonstration of model kinship is empirical. A more rigorous theoretical framework, such as the assumptions outlined in Appendix 15, is needed to explain model evolution and kinship in greater depth. e) While model kinship currently guides merging and improves performance, it does not support sustained evolution. Future progress may require incorporating environmental feedback, reward models (Silver et al., 2021), and novel architectures.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62576307), Ningbo Natural Science Foundation (2024J020), the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00), the Academic Research Fund Tier 1 (FY2025) (Grant T1 251RES2507), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ A.1 Selecting the Evolution Path
  • Performance and Kinship Evaluation: These branches were analyzed for changes in merging performance, particularly focusing on shifts in multi-task capabilities and model kinship metrics.
@@ D.1 Analysis of Model Kinship Change across Merging Stages
To determine whether the discovery of increasing model kinship in model evolution paths can be generalized to the entire model evolution process, we perform an analysis of the merging stages. Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of models at each stage. For this analysis, we randomly select 5 models from each merging stage, as delineated by boundaries identified in prior analysis - Saturation Stage ($\geq 0.75$), Improving Stage (<<0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
@@ Appendix E Analysis between Task Relatedness and Model Kinship
In the formulation of model kinship, we use the placeholder $\sim(\cdot,\cdot)$ as a similarity metric function to explore options that can effectively capture task-related differences. One such metric is cosine similarity, derived from the analysis in the task vector, which has been validated as effective for representing differences in single-task models through the cosine similarity of delta parameters (task vectors). In addition to cosine similarity, we also investigate the Pearson correlation coefficient and Euclidean distance.
However, we have not thoroughly evaluated the applicability of these metrics in the context of model evolution, particularly for merged models with multi-task capabilities. To address this, we examine the relationship between the similarity metrics and task information in subsequent sections.
@@ Appendix F Optimization Assumption of Model Evolution
Figure 9: An intuitive illustration of how model evolution can fall into local optima due to a performance-prior strategy. It shows that Merged Model 2 may be overlooked, despite its potential for better multi-task performance.
@@
Figure 11: An intuitive comparison between selective breeding and iterative model merging. The left process demonstrates breeding a tall and frutful plant by selecting parents with the desired traits in an biological scenario. The right process shows developing a model with capabilities of coding and math through model evolution.